OpenStudy (anonymous):

What is p% of 200p^3? xp^y Find x and y (index) Can you please show me the proccess? Thanks!

OpenStudy (anonymous):

\[\frac{ p }{ xp^y } (100) \] \[\frac{ 100 }{ xp^(y-1) }\]

OpenStudy (anonymous):

so \[\frac{ 100 }{ 200(p)^2 }\] percent

OpenStudy (anonymous):

which is 0.5=x and -2=y. somebody come and check this please i'm not a 100% sure. @Gravity_Dreams

OpenStudy (anonymous):

I'm really sorry, I don't understand your solution.

OpenStudy (anonymous):

sorry it was my fault, i didn't fully understand the question, just give me a second.

OpenStudy (anonymous):

\[\frac{ p }{ 100 }(200p^3)\] \[2p^4\]

OpenStudy (anonymous):

this because we divide what ever x is by 100 and add one to y, so you'll get x=2 and y=4 in this case

OpenStudy (anonymous):

Sorry, I don't really understand your solution still ^^;;

OpenStudy (anonymous):

okay so you know how to get for example f% of a number g?

OpenStudy (anonymous):

Let me dig my memory... To find f% of g.... f/100 * g right?

OpenStudy (anonymous):

exactly! so now instead of f it's p and g=200p^3

OpenStudy (anonymous):

Oh! I get it. I got up to: 200p^4/100. What do I do now?

OpenStudy (anonymous):

great, now just cancel out 200 and hundred.

OpenStudy (anonymous):

Wait, I got it. I understand now :) 2p^4. Thank you so much!

OpenStudy (anonymous):

you'll get two, it's just a simple division.

OpenStudy (anonymous):

you're welcome!